cos(180)

cos(180)

Grade:12

12 Answers

SAI SARDAR
8 years ago
-1
SEELA YUVARAJU
8 years ago
-1
SHAIK HAFEEZUL KAREEM
8 years ago
COS180=COS(180-0)=-COS1=-1 SINCE 180-0 BELONGS TO THE THIRD QUADRANT IN THIRD QUADRANT COS IS NEGATIVE SO THE VALUE OF COS180=-1
PASUPULETI GURU MAHESH
8 years ago
cos(1800) =cos (1800-00)
                =-cos (00)   since  second quadrant cos is -ve, in     1800                                     =   cos=cos,sin=sin  ther   is  no change             
              =-1
RIYAZSHAIK
8 years ago
-1 because cos(180)=cos(180+0)=-cos(0)=-1 becase the angle lies on 2nd quadrant thats why the cos sign is – so the answer is -1
mgayathri
8 years ago
-1
cos value decreases as the angle increases so cos 0=1
slowly the value decreses resulting in cos180=-1
vamsi krishna
8 years ago
cos(180+x) is in 3rd quadrant.in 3rd quadrant cos0 is negative.then cos0 is 1.......-cos0=-1.....here cos(180) degrees is forms that cos(180+0) i.e., -cos(0)=-1...............in the 4 quadrants you should remember that one formula “All Silver Tea Cups” that means ALL--all trigonometric ratios are positive...................S means sine & cosec are positive remaining are negative................T means Tan & cot are positive & remaining are negative............C means cos & sec are positive & remaining are negative................
VENKI
8 years ago
cos180=-1 cos(180-0)=-cos0   [cos=1] so -cos0=-1  because 180 is second (2) quadrant so cos180=-1
AYESHA
8 years ago
cos(180)=-1.cos is in 4th quadrant. 180 angle lies in 2nd quadrant. So cos become negative here. So the solution is in negative form.
D jaya
8 years ago
cos(180)=-1 cos is forth quadrent. so cos is negative.  the angle 180 is secand quadrent. then cos 0 is 1 it is possitive.
raj
8 years ago
hello,
cos(180)=-1....
karthikeya
7 years ago
cos(180)=-1...